Weekends

Given two dates D1 and D2, write a program to count the number of Saturdays and Sundays from D1 to D2 (including D1 and D2).

The date in string format is like "8 Feb 2021".Input


The first line of input will contain date D1 in the string format.

The second line of input will contain date D2 in the string format.


Output

The output should be a single line containing two integers separated by space


Explanation

For example, if the given dates are "25 Jan 2021" and "14 Feb 2021", the Saturdays and Sundays dates from "25 Jan 2021" to "14 Feb 2021" are

"30 Jan 2021" is a Saturday

"31 Jan 2021" is a Sunday

"6 Feb 2021" is a Saturday

"7 Feb 2021" is a Sunday

"13 Feb 2021" is a Saturday

"14 Feb 2021" is a Sunday


output

Saturday: 3

Sunday: 3


Sample Input 1

25 Jan 2021

14 Feb 2021


Sample Output 1

Saturday: 3

Sunday: 3


Sample Input 2

25 May 2019

22 Dec 2021


Sample Output 2

Saturday: 135

Sunday: 135


1
Expert's answer
2021-03-28T03:36:38-0400
from datetime import date


def numOfDays(D1, D2):
    num=(D2 - D1).days
    if(num%7==0):
     print('Number of saturday and sunday:',int(num/7))
    else:
     if(num%7<5):
      print('Number of saturday and sunday:',int(num/7))
     else:
      if(num%7<6):
       print('Number of saturday:',int(num/7)+1,'and number of sunday',int(num/7))
    return num



# Driver program
D1 = date(2021, 3, 10)
D2 = date(2021, 3, 2)
print('Number of days between these two dates',numOfDays(D2, D1))